### v1

Abstract
Fast and accurate depth sensing has long been a significant research challenge. Event camera, as a device that quickly responds to intensity changes, provides a new solution for structured light (SL) systems. In this paper, we introduce Gray code into event-based SL systems for the first time. Our setup includes an event camera and Digital Light Processing (DLP) projector, enabling depth estimation through high-speed projection and decoding of Gray code patterns. By employing spatio-temporal encoding for point matching, our method is immune to timestamp noise, realizing high-speed depth estimation without loss of accuracy. The binary nature of events and Gray code minimizes data redundancy, enabling us to fully utilize sensor bandwidth at 100%. Experimental results show that our approach achieves accuracy comparable to state-of-the-art scanning methods while surpassing them in data acquisition speed (up to 41 times improvement) without sacrificing accuracy. Our proposed approach offers a highly promising solution for ultra-fast, real-time, and high-precision dense depth estimation. Code and dataset will be publicly available.
Keywords:
Structured light Depth estimation Event camera Dynamic vision sensor
1 Introduction
Depth estimation plays a pivotal role in various fields[17, 42, 7], including machine vision and robotics, with applications spanning 3D modeling, autonomous driving, industrial inspection, and more. As a depth estimation method based on the principle of triangulation, the structured light system (SL system) projects a known pattern onto the target and estimates depth based on the pattern spatially-modulated by the target[2, 36]. The SL system is favored for its high accuracy, dense depth estimation capability, insensitivity to target texture features, and reasonable cost, rendering it indispensable in domains such as 3D scanning and augmented reality[45]. Presently, the SL system boasts numerous mature commercial applications, underscoring its widespread adoption and effectiveness.
However, SL systems employing traditional frame-based cameras are restricted to low frequency (usually at a frame rate of tens of hertz) due to bandwidth and exposure time constraints of the cameras. While high-speed cameras offer the potential for ultra-fast acquisition[21], their exorbitant cost renders them impractical for many applications. Moreover, traditional SL systems suffer from data redundancy. Images captured by frame-based cameras contain substantial redundant information, particularly in point and line-structured light systems [1]. This redundancy undermines the acquisition speed and overall efficiency of the system.
In recent years, event cameras[3, 10, 24], also known as Dynamic Vision Sensors (DVS), have emerged as a promising option for SL systems. These bio-inspired devices asynchronously report changes in light intensity at the $\mu$s level. Due to their capability of solely generating signals for changes in light intensity, event cameras inherently mitigate data redundancy in most SL systems at the hardware level. Additionally, the high dynamic range (HDR) of event cameras allows them to operate effectively under diverse illumination conditions. Consequently, event cameras offer an alternative solution for high-speed depth estimation in SL systems.
Some works have already explored its potential in depth estimation. Laser-point projector based approaches[26, 30, 28] have achieved a scanning speed of 60Hz and high recovery accuracy. After program engineering, they can perform real-time dense target depth estimation at lower frame rates. However, this method is limited by the low encoding efficiency and timestamp noise (Latency and Jitter) , making it hard to achieve higher-speed depth estimation. On the other hand, works based on the full-frame method[23, 25, 40] can achieve higher speed by using Digital Light Processing (DLP) projector. However, these methods are still affected by timestamp noise and face difficulties in robust decoding.
Therefore, considering the high-speed response and low redundancy characteristics of event cameras, as well as the high-speed projection characteristics of DLP projectors, we propose our Structured light system based on Gray code with an event camera (SGE) , which uses Gray code to encode the target, and estimates its depth based on epipolar geometry. This method first utilizes spatio-temporal encoding instead of timestamp matching for point matching, allowing for faster DLP projection and not being affected by noise of event camera and projector. Secondly, the binary nature of Gray code perfectly match with the binary response of event cameras. This allows our method to preserve all encoding information without data redundancy. Thirdly, we have developed a novel calibration scheme for event-based SL systems. Compared to previous schemes[27, 39, 9], our scheme has sub-pixel level accuracy and is convenient to use, and it also applicable to all event-based SL systems. Finally, we also propose the Gray code X-map (GX-map) disparity query scheme. This scheme enable us to directly query disparities instead of disparity searching. We achieved a processing speed improvement of three orders of magnitude in our experiment. These advantages allow us to achieve a much higher depth estimation speed exceeding 1000 Hz while maintaining comparable accuracy to state-of-the-art methods at a millimeter-level precision, and within a range of 1/200 of the average depth.
The contributions of this paper are as follows:
  • Structured light system based on Gray code with an event camera system (SGE) is proposed with depth acquisition speed over 1000 Hz (Up to 41 times improvement). Experiments on both static and high-speed dynamic scenes show the superiority of the proposed SGE.
  • A sub-pixel level, universally applicable calibration method and the GX-map disparity query scheme are proposed. The latter improves the processing speed by a factor of 1000.
[…]
2 Related work
Structured light system. Traditional SL system performs depth estimation based on patterns spatial modulated by the target. As a field that has been researched for many years, it has yielded many accomplishments. Such as different projection pattern: point[1, 6, 14], line[5], patterned area[29, 49], binary coding[19, 16, 44], continuous coding[22, 15]. And as well as different numbers of pattern, including multiple-shot[47] and single-shot[38, 20, 48]. These schemes have their own advantages and disadvantages in terms of accuracy, speed, environmental adaptability, etc., thus requiring selection based on specific circumstances[13, 46]. However, due to the limited frame rate of frame-based cameras, typically only tens of Hz, it is impossible for traditional SL systems to achieve high-speed depth estimation. This limitation prevents their application in dynamic scenes.
Event-based depth estimation. Due to advantages[10] of event cameras, for example high temporal resolution, they have been gradually used in the field of depth estimation. Initial approaches to event-based depth estimation primarily relied on feature matching with multi-view stereo techniques [32]. Subsequently, event-based Simultaneous Localization and Mapping (SLAM) methods [33, 35, 50] gained traction for constructing environmental representations or maps and deriving metric depth information. However, these approaches rely on either the scale provided by available camera poses or auxiliary sensors like inertial measurement units (IMUs) for metric depth recovery. Gallego et al. [11] removed dependencies on other modalities with a contrast maximization framework. Later, learning-based methods utilized events alone to predict depth with multi-view and egomotion constraints. Ye et al. [43] proposed a self-supervised framework for depth, optical flow, and egomotion prediction based on the principle of the egomotion model, which was further improved by Zhu et al. [51] with Temporal Reprojection Loss, Stereo Disparity Loss, and other constraints. These methods all utilized multi-view approaches, either stereo cameras or egomotion. Hidalgo et al. [18] were the first to propose predicting dense depth with a monocular camera using supervised learning. Gehrig et al. [12] combined frames and events for more robust depth estimation. However, learning-based methods still have significant limitations in terms of the accuracy of depth estimation. This heavily rely on high-quality datasets and proper training techniques. Furthermore, since the data is generated from slow target motion and need to pass through a slow network, both the data collection and processing speed are relatively slow.
Event-based SL system. Event cameras have also gained attention in SL systems. Early work[4] used a 500Hz pulsed laser to achieve three-dimensional reconstruction by moving the target. However, this method required a displacement stage for scanning. Matsuda etal.[26] proposed Motion Contrast 3D (MC3D) replacing the pulsed laser with a laser-point projector and utilizing pixel timestamps for point matching. This approach laid a foundation for subsequent research but suffered from quality degradation under high-speed scanning due to timestamp noise. Muglikar etal.[30] analyzed this noise and improved MC3D’s matching algorithm, enhancing robustness against timestamp noise at a scanning speed of 60Hz. However, it slowed down processing speed due to increased computational effort. Morgenstern etal.[28] introduced X-map method, which used table look-up operations for fast disparity calculation and achieved real-time depth estimation. However, timestamp noise interference remained a challenge. Other works [23, 25, 40] explored DLP projection for faster depth estimation but faced issues of realizing high accuracy and robust decoding. In summary, most existing works struggle to achieve high-accuracy, high-speed, and robust depth estimation simultaneously, but showing favorable results only in specific aspects.
3 Method
We first introduce the foundation and basic consideration of our SGE method in Sec. 3.1 and then outline the pipeline in Sec. 3.2. In Sec. 3.3, our simple and high accuracy calibration scheme for event-based SL system will be proposed. Finally, we detail the depth estimation scheme based on GX-map query and Time-overlapping strategy in Sec. 3.4.
3.1 Preliminaries and Analysis
SL system for depth estimation. Traditional SL systems, composed of a frame-based camera and a projector, typically employ triangulation for depth estimation. Similar to common stereo vision system, SL systems can be conceptualized as artificially projecting "features" onto the target surface using a projector. Due to the typically dense nature of this projection, SL systems can achieve dense depth estimation that surpasses the capabilities of stereo vision alone, particularly in generating highly accurate depth information at short distances [37]. The primary distinction among SL schemes lies in the use of different projection patterns. But these patterns only lead to variations in depth encoding and do not impact the disparity calculation process. Therefore, in structured light depth estimation, after obtaining the depth encoding, stereo rectification is first performed based on the calibration parameters, and then point matching:
$P_{pr}(x_{pr},y_{pr})=F(P_{cr}(x_{cr},y_{cr})),$ (1)
where $P_{cr}$ denotes an encoded point in the camera’s pixel coordinate system, $P_{pr}$ represents the corresponding point in the projector’s pixel coordinate system. $x_{pr}$, $y_{pr}$, $x_{cr}$, $y_{cr}$ represent pixel coordinates in the corresponding pixel coordinate system. And $F$ represents the matching algorithm. Assuming that the disparity direction is along the $X$ direction, the depth can be obtained by:
$Z(P)=\frac{f\cdot b}{x_{cr}-x_{pr}}.$ (2)
Here, $f$ denotes the focal length, while $b$ represents the baseline distance. However, SL systems encounter two basic limitations preventing it from high-speed measurement. First, it is restricted by low frame rate of the frame-based camera, typically only tens of hertz. Second, the abundant data redundancy of images not only takes up sensor bandwidth but also demands additional data processing[30].
Event generation. The recently popular event cameras provide a novel solution for high-speed structured light depth estimation thanks to their high temporal resolution. Each pixel of the event camera reports the intensity changes after perceiving intensity change exceeding the threshold, which we call events ($E=(x,y,\tau,p)$), where $x$, $y$, $\tau$, $p$ represent the pixel coordinates, output timestamp, and polarity, respectively. For an event $e$, it’s polarity $p_{e,\tau}$ can be formulated as:
$p_{e,\tau}=\begin{cases}+1,\text{if}\log\left(\frac{\mathcal{I}_{t,e}}{\mathcal{I}_{(t-\Delta t),e}}\right)>C_{positive},\\ -1,\text{if}\log\left(\frac{\mathcal{I}_{t,e}}{\mathcal{I}_{(t-\Delta t),e}}\right)<C_{negative},\end{cases}\tau=t+\sigma$ (3)
where $\mathcal{I}$ represents intensity, $C_{positive}$, $C_{positive}$ represent the threshold for generating positive and negative events, which can be manually adjusted. $t$, $\sigma$, and $\tau$ are the ideal timestamp, the timestamp noise, and the actual output timestamp, respectively.
The previous works[30, 28] have utilized event-based SL system to realize depth estimation. However, these point scanning approaches suffer from timestamp noise and low encoding efficiency. The former leads to a decrease in accuracy when scanning speed increases. The latter means more timestamps are needed for depth encoding, as shown in Equ. 4, where N is the number of timestamps needed for encoding, C is the number of columns (disparity direction).
$N=C,\vspace{-1.5em}$ (4)
$N=\log_{2}(C),$ (5)
Therefore, for the first time we propose using Gray code as depth encoding pattern for depth estimation in event-based SL systems. Firstly, by using Gray code, we achieve higher encoding efficiency while maintaining the absence of redundant information, as shown in Equ. 5. Secondly, our approach is immune to timestamp noise because we only need timestamps for data segmentation. Therefore, our SGE has the capacity to higher-speed and high-precision depth estimation.
Figure 1: The pipeline of the proposed SGE. The overall process starts with system calibration, and obtains depth after data acquisition and data processing.
3.2 Overall Pipeline
Our SGE comprises a DLP projector and an event camera, as depicted in Fig. 1. Before deployment, we first calibrate the system (Sec. 3.3) using our proposed Simple Event-based SL system Calibration (SEC). Then the projector independently and continuously projects Gray code onto the target. Thanks to DLP’s expertise in binary projection, our system can achieve an exceptionally rapid projection speed of up to 2487HZ, which is only limited by our hardware. Patterns modulated by the target depth are captured by the event camera from another view. As the data from the event camera is in the form of an event stream, segmentation according to each projection and subsequent conversion into image form are required (Step (a)). After acquiring a whole set of patterns, we use the traditional Gray code algorithms[31] to derive the depth-encoded image (Step (b)). Finally, leveraging the system’s calibrated parameters (Step (c)) and GX-map (Step (d)) , we determine the depth in Sec. 4. For dynamic scenes, we propose the Time-overlapping strategy to achieve higher data utilization and depth estimation speed (Sec. 5). Alg. 1 provides the pseudo-code of the proposed SGE.
Figure 2: Pipeline of calibration for event-based SL system. Event data from different angle are needed for calibration.
3.3 Simple Event-based SL System Calibration
Figure 3: The calibration reprojection error of the projector and camera. The horizontal and vertical axes are measured in pixels.
Previous work[30] converts the event to images with E2VID [34] and utilize calibration tool [27] for rough extrinsic parameters calibration. Alternatively, some[28] adopt a step-by-step strategy and separately calibrate the camera’s intrinsic parameters. However, these methods (i) necessitates dual calibration procedures and complex calibration process, rendering it cumbersome to use. And (ii) images generated using E2VID may struggle to maintain sub-pixel level authenticity, while the step-by-step strategy introduces significant artificial accuracy loss. While some works have achieved high-precision calibration [39], they are limited to laser-point projectors or event cameras with gray mode.
Therefore, based on Gabriel’s calibration tool[8], we propose a novel, sub-pixel level, simple, and versatile event-based SL system calibration scheme SEC, as shown in the Fig. 2. Initially, we affix an LCD screen presenting a flickering chessboard onto a planar surface and project another chessboard pattern onto the other side. Following the acquisition of multiple datasets from various perspectives, the data is directly binarized and subjected to calibration procedures using the standard calibration tool. Our approach facilitates simultaneous calibration of all parameters through straightforward operations, significantly mitigating the complexity associated with calibration processes. Furthermore, our method attains high calibration precision, demonstrating the capability to achieve sub-pixel level reprojection errors. As evidenced in Fig. 3, we achieve the average reprojection error of [0.59, 0.54] for the projector and [0.40, 0.40] for the event camera in the experiment with 23 sets of data collected from different poses. Moreover, our SEC is universally applicable to all event-based SL systems.
Input : Depth-encoded image of projector: $D_{p}$
Event data: $E$
Calibration(intrinsic and extrinsic parameters): $C$
Output : Depth map: $D$
/* Initialization */
// Obtain the rectified depth-encoded image of projector
$D_{pr}$leftarrow ($D_{p}$, $C$) ;
// Obtain GX-map according to Equ. 6
$GX()$leftarrow $D_{pr}$ ;
// Time-overlapping strategy
/* Main loop */
// Event data is not over
while E $\neq$ None do
       // Sequential acquisition of Gray code slice data and binarization (Step (a))
       $G_{x}$leftarrow $E$
       // A complete set of Gray code is obtained
       if $n(G)$ $=$ $N$  then
             // Depth-encoded image of camera calculation (Step (b))
             $D_{c}$leftarrow $G$
             // Rectification (Step (c))
             $D_{cr}$leftarrow ($D_{c}$, $C$)
             // Disparity query according to Equ. 7 (Step (d))
             $dis$leftarrow ($D_{cr}$,$GX()$)
             // Depth map calculation according to Equ. 2
             $D$leftarrow ($dis$, $C$)
       end if
end while
Algorithm 1 Depth estimation by SGE method
3.4 Depth Estimation
Depth-encoded Image Acquisition. As mentioned in Sec. 3.2, the continuous projection data captured by the event camera manifests in the form of an event stream. Consequently, the initial step involves segmenting the data based on each projection. The presence of projection dark time intervals, as illustrated in Fig. 4, enables straightforward identification of time windows within the event stream devoid of event occurrences, which serve as natural boundaries for segmentation. Subsequently, the segmented slices are binarized into Gray code images, following which traditional Gray-coded depth estimation algorithms [31] are employed to derive depth-encoded image of the target, as demonstrated in Fig. 5.
Figure 4: The projection characteristics of DLP projectors and laser-point projectors. (a) and (c) show the timestamp and time gradient variations with event id (sorted by time). (b) and (d) present the scatter plot of the events during the projection. DLP employs multi-point scanning, while laser-point uses raster scanning with non-linear characteristics. Both include dark intervals between projections..
GX-map Disparity Query. Following the acquisition of depth-encoded images, point matching is essential to extract depth information. Due to the high time-consuming nature of epipolar line search, we propose the GX-map disparity query scheme tailored for Gray code. Upon conducting stereo rectication to obtain rectified depth-encoded images, we store the x coordinates $x_{pr}$ of the projector’s rectified depth-encoded image over y coordinates $y_{pr}$ and the encoding value $v_{pr}$ as GX-map, then directly query the corresponding disparity $d$ using the rectified camera’s depth-encoded image. This is formulated as:
$v_{pr}=RP(x_{pr},y_{pr})\rightarrow x_{pr}=GX(y_{pr},v_{pr})$ (6)
$d=x_{cr}-GX(y_{cr},v_{cr})$ (7)
where $x_{cr}$, $y_{cr}$ and $v_{cr}$ represent the coordinates and encoding value of camera’s rectified depth-encoded image, respectively. $RP()$ and $GX()$ represent function that queries based on projector’s rectified depth-encoded image and GX-map. Given that the rectified depth-encoded image of the projector remains unchanged under identical intrinsic and extrinsic parameters, the GX-map generation operation requires execution only once and can subsequently facilitate multiple queries. Through GX-map querying, we have attained a thousand-fold enhancement in computational speed compared to traditional epipolar line search on the same processor, making high-speed real-time depth estimation possible.
Figure 5: Depth-encoded images. Different colors represent different depth encodings. The red box indicates the approximate matching position.
Time-overlapping. For dynamic scenes, due to the fact that a set of Gray codes starting with any serial number can compose a depth-encoded image, therefore, similar to[41], we propose a Time-overlapping strategy to maximize data utilization, as illustrated in Fig. 6.
Figure 6: Time-overlapping strategy of Gray Code. We take a 4-bit Gray code as an example, each consecutive 4 images can generate a corresponding depth-encoded image, greatly improving data utilization and depth estimation speed.
For depth estimation using N-bit Gray code, N projection patterns is required to compute depth. After capturing the initial set of binarized Gray codes, each subsequent Gray code can be combined with the previous images to calculate a new depth-encoded image. This means each projection corresponds to a new depth estimation, further improving the speed of the depth estimation by N times for N-bit Gray code.
4 Experiments
4.1 Settings
[…]
Figure 7: The hardware setup, involving two projectors with different field of view (FoV) and resolutions. Calibration factors are taken into account, resulting in differences in the structure of the two setups.
In the experiment of static scenes depth estimation, we use the Sony Mobile projector MP-CL1A with a resolution of 1920$\times$720 pixels and a fixed projection speed of 60Hz. It is a laser-point projector working in a raster scanning pattern, which projects clear images at any distance. However, it suffers from laser point diffusion while projection.
[…]
Compared methods. We choose the state-of-the-art point scanning methods: MC3D[26], the initial and optimized ESL[30], and the X-map[28] in the scanning speed of 60HZ as our comparing methods. We average the result from MC3D[26] within one second following [30]. It is worth mentioning that our calibrated parameters are adopted by other methods.
In this task the ground truth is difficult to obtain because it is hard to find a method can generate dense depth of natural scenes above millimeter level accuracy. Therefore, ESL[30] uses averaged MC3D as their ground truth, while X-maps[28] uses optimized ESL. Others[25, 23] choose qualitative comparison only. Although we also think the averaging operation can improve the accuracy of depth estimation, which means averaged MC3D may be the most accurate method among others, we think because they both using the same point scanning method and vary only in different data processing process, therefore the comparison between them may also reflect the similarity of their depth results. This means the accuracy may be over-evaluated by selecting the same type of method as the baseline. Therefore it is unfair and impossible for our SGE to surpass other point scanning methods if we choose a point scanning method as ground truth.
But because our SGE uses different types of data, comparing our method with point scanning methods can fairly reflect our absolute precision. Therefore, we do not set ground truth in our comparison but compare our method with each point scanning method to verify the absolute accuracy of our scheme.
Metrics. The root mean square error (RMSE) and the fill-rate (FR) are adopted as our evaluation metrics. To minimize the influence of data differences, like slight differences in edge positions, we only use solid regions in both methods to calculate RMSE. We define solid regions as where both methods give a reliable depth estimation, which means the difference between them is within the threshold. The FR measures the ratio of points below depth error threshold in SGE to measured points in compared method. The depth error threshold is 1% of the average scene depth.
4.2 Static Scenes
Due to the inability of the point scanning schemes to utilize DLP projectors, we adopt the laser-point based system (as shown in Fig. 7 (b)) to eliminate device differences. In our experiment, we captured multiple scenes with varying complexity, including highly detailed gypsum statues, surfaces with fewer details, and everyday scenes to show the versatility of our solution. To minimize variables, we keep the projection light intensity, ambient light intensity, camera threshold, and shooting position constant while capturing two types of data. During the experiment, we use mean filtering to reduce discontinuity. Since the projector’s 720 horizontal resolution cannot be fully utilized by Gray code encoding, we use only 512 columns for projection, which results in the missing background on the right side.
Figure 8: Qualitative results of different methods on static scenes with the laser-point based system. Our SGE not only shows the performance consistent with the the-state-of-art schemes, but also presents more solid results in concave regions and less noise in overexposed and underexposed locations.
Table 1: Quantitative results of the differences between our SGE method and other methods in static scenes. RMSE and FR are adopted as evaluation protocol.
Scene David Dragon Structure gslim boxes fan bottle wizard arena column
Mean Depth 450mm 410mm 530mm 360mm 510mm 530mm 320mm 400mm 430mm 320mm
[…]
MC3D-$1s$ 0.93 1.09 0.99 0.95 0.97 1.58 0.95 1.43 0.98 1.76 0.99 1.64 0.98 0.73 0.97 1.01 0.95 1.09 0.97 0.76
ESL 0.95 1.00 0.99 0.76 0.94 2.0 0.96 1.17 0.98 1.46 0.99 1.28 0.99 0.62 0.98 0.80 0.96 1.04 0.98 0.64
ESL-opt 0.95 0.99 0.99 0.90 0.97 1.66 0.95 1.40 0.98 1.70 0.99 1.62 0.99 0.68 0.98 1.00 0.96 1.12 0.99 0.72
Xmap 0.94 1.22 0.99 0.81 0.92 2.33 0.94 1.11 0.97 1.43 0.97 1.56 0.98 0.66 0.97 0.80 0.96 1.19 0.98 0.65
Fig. 8 presents the qualitative depth estimation results. It can be observed that for point scanning methods, due to timestamp noise and laser point diffusion, there is more error present at the concave regions (e.g. Structure, Boxes) and overexposed and underexposed locations (e.g. Wizard, Arena, Column). This is more pronounced in the X-map scheme, which does not utilize interpolation or averaging methods, and slightly improved in the ESL scheme with post-processing and in the MC3D scheme by averaging. However, our SGE demonstrates strong resistance to timestamp noise and laser point diffusion and performs accurate depth estimation in any case.
The quantitative results are shown in Tab. 1. Our method achieves a RMSE of less than 2mm, which is less than 1/200 of the average depth, and a FR over 0.93 when compared to other state-of-the-art methods. This indicates that our method achieves high accuracy comparable to the state-of-the-art with the absolute precision at the millimeter level. During the comparison, considering the potential application of real-time depth estimation, we just applied simple and fast median filtering in the pre-processing and post-processing stages to optimize the results. It is worth mentioning that our method can be further improved if additional off-line optimization algorithms are adopted, like those used in ESL.
4.3 Results Under Different Scanning Speed
[…]
4.4 Dynamic Scenes
[…]
Because obtaining recognized high-precision depths of natural scenes at a $\mu s$ resolution is difficult, if not impossible, without employing methods that rely on high-speed cameras or other costly equipment. Therefore, we cannot prove the accuracy of SGE directly. But as the experiment shown in Sec. 4.3 proves the high accuracy of our SGE at high scanning speeds, it can be assured that the accuracy of our scheme in dynamic scenes is guaranteed. The only additional factor impacting the accuracy is the incorrect depth encoding caused by motion, as the noise around the edge shown in Fig. 10. This is a problem that every dynamic scene depth estimation method will encounter. The only solution is to enhance data acquisition, which is the greatest strength of our SGE.
Figure 10: Visual results of our SGE on dynamic scenes. Various dynamic scenes at diverse timings are presented, proving our capacity of depth estimation on dynamic scenes
5 Conclusion
In this work, We propose a high-speed event-based structured light depth estimation scheme: SGE. For the first time we introduce Gray code into event-based SL systems for depth estimation with a readily available DLP projector, realizing higher encoding efficiency, timestamp noise immunity and higher depth estimation speed. In addition, due to the absence of a simple and reliable calibration scheme, we present a sub-pixel level, universally applicable calibration method SEC for event-based SL system. Furthermore, we introduce the GX-map disparity query scheme and the Time-overlapping strategy to enhance computational speed and data utilization. We also captured a dataset consists of static and dynamic scenes with DLP and laser-point based system. Corresponding calibration data is also provided. Experimental results show that our SGE achieves kilohertz-level data acquisition speed while maintaining comparable accuracy as the state-of-the-art point scanning methods. This proves the capability of our SGE to realize high-speed and high-accuracy depth estimation.
In the future, we insist that our SGE can still be promoted by using more advanced event cameras with shorter refractory period and projectors with higher scanning speed. In addition, we believe a calibration scheme with a higher field of view (FoV) utilization and the study in reducing decoding errors caused by motion through data processing can also enhance our method.